### 26 Feb 2024 v1

@@ Appendix A Environment models
We use a standard deep learning workflow to train the environment models. The different optimization techniques are described in detail by Raff [27]. Table 2 provides an overview of the most important hyperparameters. Since the amount of training data is fairly small, we employ 8 CPU cores to train a single model. Note that the models predict the individual force coefficients rather than the full reward. Consequently, the output layer has 12+2 and 14+6 neurons for the cylinder and pinball test cases, respectively. The training is stopped once the maximum number of epochs is reached, the absolute loss value falls below a threshold value, or the relative change of the loss averaged over 40 epochs becomes too small.

### 10 Apr 2024 v2

@@ Appendix A Environment models
We use a standard deep learning workflow to train the environment models. The different optimization techniques are described in detail by Raff [31]. The activation function, the number of hidden layers, and the number of neurons per layer have been determined by employing a simple grid search. Table 2 provides an overview of the most important hyperparameters. Since the amount of training data is fairly small, we employ 8 CPU cores to train a single model. Note that the models predict the individual force coefficients rather than the full reward. Consequently, the output layer has 12+2 and 14+6 neurons for the cylinder and pinball test cases, respectively. The training is stopped once the maximum number of epochs (gradient descent steps) is reached, the absolute loss value falls below a threshold value, or the relative change of the loss averaged over 40 epochs becomes too small.
@@ Appendix B PPO hyperparameters
Appendix C List of abbreviations
AFC active flow control
CFD computational fluid dynamics
DRL deep reinforcement learning
IDDES improved delayed detached eddy simulation
MB model-based
MEPPO model ensemble proximal policy optimization
METRPO model ensemble trust region policy optimization
MF model-free
RL reinforcement learning
Table 4: Abbreviations used throughout the article.